# Telegrapher’s Generative Model via Kac Flows

### v1

@@ 5.2 Multi-Dimensional Kac Process
$\frac{C_{1}}{(8\pi)^{\frac{d}{4}}}\left(\frac{1}{at^{1+\frac{d}{4}}}+\frac{1}{ca^{\frac{1}{2}}t^{\frac{1}{2}+\frac{d}{4}}}\right).$ (80)
@@
$\leq\Big{(}C_{1}\Big{(}\frac{1}{at}+\frac{1}{c(ta)^{\frac{1}{2}}}\Big{)}+\frac{1}{(8\pi t)^{\frac{1}{4}}}\Big{)}$ (83)
@@ 6 Velocity Field of the Multi-Dimensional Kac Flow
In this section, we want to develop a method how to retrieve the velocity field $v_{t}$ of the Kac flow in multiple dimensions. By conditioning on single points $x_{0}\in\mathbb{R}^{d}$, the multi-dimensional velocity field $v_{t}(x|x_{0})$ decomposes into univariate velocities $v_{t}^{i}(x^{i}|x_{0})$ which are analytically and numerically accessible. These conditional velocity fields can then be used to approximate the true velocity field $v_{t}$ within the framework of conditional flow matching. But note that conceptually, our velocities are conditioned to start at $t=0$ in the target measure – reversing the direction of usual flow matching.
@@
Furthermore, assume that the one-dimensional measure flows $\mu^{i}_{t}$ satisfy a continuity equation with some vector field $v^{i}_{t}$, i.e.
@@
By Proposition 11, $\mu_{t}(\cdot|x_{0})$ admits a velocity field $v_{t}(\cdot|x_{0})$ satisfying the continuity equation (2). By construction, the components $(X_{t}^{1},...,X_{t}^{d})$ of $\mathbf{X}_{t}$ are mutually independent. Hence, $\mu_{t}(x|x_{0})=\prod_{i=1}^{d}\mu_{t}^{i}(x^{i}|x_{0}^{i})$, where the one-dimensional probability flows $\mu_{t}^{i}(\cdot|x_{0}^{i})$ fulfill a continuity equation with 1D velocity fields $v_{t}^{i}(\cdot|x_{0}^{i})$. By Lemma 13, it holds the decomposition (88), and Theorem 4 immediately gives the formula (93). Finally, [17, Theorem 3.3], resp. [36, Theorem 6.3] provides (94). ∎
Having identified the conditional vector field $v_{t}(\cdot|x_{0})$ for the conditional distribution $\mu_{t}(\cdot|x_{0})$ of the Kac process, we get the following result by employing [15], see also [36, Theorem 5.1, Remark 5.3], with slight adjustments to our setting.
@@
but we have no direct access to $v$. However, following the same lines as in [15], we see that
@@
with a constant not depending on $\theta$ and the Conditional Flow Matching (CFM) loss is given by
@@
Remark 17 (Mean Reverting Kac Flow).
In general, the ’standard’ Kac process
$X_{t}:=X_{0}+K_{t}$
and $K_{t}$ themselves are not equal at the final time $T$. It is common practice in generative modeling to augment the process such that the data does no longer contribute to the final distribution (compare with variance preserving diffusion processes, e.g. the Ornstein-Uhlenbeck process). In the same fashion, we introduce the ’mean reverting’ Kac process as
$M_{t}\coloneqq f(t)X_{0}+K_{t},\,t\in[0,T],$
for some smooth $f$ satisfying $f(0)=1$ and $f(T)=0$. In our numerical examples we will use $f(t)=1-t/T$. Now we have at time $T$ that $M_{T}=K_{T}$. Furthermore, it holds
$\dot{M}_{t}=\dot{f}(t)X_{0}+\dot{K}_{t}$
and the conditional velocity of the process $M_{t}$ is given by
$v_{M}(t,x|x_{0})$ $=\mathbb{E}[\dot{M}_{t}|M_{t}=x,X_{0}=x_{0}]=\mathbb{E}[\dot{f}(t)x_{0}+\dot{K}_{t}|K_{t}=x-f(t)x_{0},\,X_{0}=x_{0}]$ (98)
$=\dot{f}(t)x_{0}+v_{K}(t,x-f(t)x_{0}|x_{0}),$ (99)
where $v_{K}(t,x|x_{0})$ is the conditional Kac velocity field (88). Using the same arguments as in Corollary 15, we can therefore also recover the vector field associated to the mean reverting Kac process.
@@ 7 Numerical Experiments
In this section, we use the flow matching framework to train a neural network to approximate the velocity field from Corollary 15 by minimizing the CFM loss given in (97). To this end, we compute the conditional velocity $v_{t}(x|x_{0})$ via its analytic formula (93), resp. (98). In order to sample from $\mathbf{X}_{t}\sim u_{t}(\cdot|x_{0})$, where $x_{0}$ is fixed, we only need to sample i.i.d. components $X_{t}^{i}$ of the Kac process starting in 00. This can be done efficiently, since we have access to the analytic form (31) of the one-dimensional Kac distribution starting in 00 (use e.g. inverse transform sampling). Further, if we wish to sample fixed realizations of a Kac process for varying times $t$, we can efficiently simulate the random walk itself as explained in Appendix A.1.
During inference, we sample the latent variable from the Kac process $K_{T}$, which starts at zero at a final time $T$, and then integrate the flow ODE (5) backward in time using an ODE solver with the learned vector field as the drift. Using the standard Kac flow introduces an approximation error, namely the distance between $X_{T}$ and $K_{T}$. This can be avoided by applying the mean reverting Kac flow from Remark 17.
7.1 Diffusion vs. Kac: A 2D Experiment with Dirac-like Modes
@@
We apply the conditional flow matching framework to both the Kac and diffusion based model. Here we use the standard Kac process starting in the data with parameters $\frac{c^{2}}{a}=1$, and the standard diffusion with parameter $\sigma=1$ and the velocity field (11).
@@
7.2 MNIST
Figure 6: The standard forward Kac process starting in MNIST data, simulated until time $T=10$ with $(a,c)=(1,1)$ (upper row) and $(a,c)=(25,5)$ (lower row).
In Figure 6 we visualize the forward Kac process. We now use the Kac process for generative modeling as described in Remark 17, i.e. the mean reverting Kac process. We train the velocity field using conditional flow matching with the vector field (95) and compare a variety of choices for $(a,c,T)$. We apply a UNET architecture to parametrize the velocity field and use an ODE solver to simulate the corresponding flow ODE. For further details we refer to the implementation details in the Apppendix A.2. Note that we fixed the random seed for each run.
In Figures 7 and 8 we plot the generation results. We see that for $(a,c)=(1,1)$ the corresponding flow is irregular, leading to worse generation results. Increasing $(a,c)$ stabilizes the model and we see consistent performance. Increasing the final time $T$ seems to slighly improve generation results for $(a,c)=(1,1)$. We additionally plot generated trajectories in the Appendix A.2.
$(a,c)=(1,1)$
$(a,c)=(25,5)$
$(a,c)=(100,10)$
$(a,c)=(225,15)$
$(a,c)=(900,30)$
Figure 7: Generation results of the mean-reverting Kac model for $T=1$.
$(a,c)=(1,1)$
$(a,c)=(25,5)$
$(a,c)=(100,10)$
$(a,c)=(225,15)$
$(a,c)=(900,30)$
Figure 8: Generation results of the mean-reverting Kac model for $T=5$.
@@ Acknowledgement.
J. Chemseddine acknowledges funding of the DFG project STE 571/17-2 within the SPP 2298 "Theoretical Foundations of Deep Learning". R. Duong, P. Friz and G. Steidl acknowledge funding of the project EF-LI-Opt-2 within the DFG Excellence Cluster MATH+. We thank Gregor Kornhardt and Nicolaj Rux for fruitful discussions.
@@ A.2 Implementation Details
i.e. we draw $X_{0}$ directly from our GMM and add the exact Kac noise $K_{t}$, thereby avoiding any sampling approximation in this toy example. All models were trained for $500{,}000$ iterations with a learning rate of $5\times 10^{-4}$ and a batch size of $256$. For validation we compute the negative log-likelihood (NLL) of $N$ model-generated samples under the true GMM density:
@@
MNIST Experiment
We trained all models for $70$k iterations, a batch size of $128$ and a learning rate of $5\times 10^{-4}$. We used the torchdiffeq [5] package to simulate the ODEs with an explicit euler scheme. Furthermore we used an exponential moving average on the network weights, with a decay factor of $0.99$. We use a lightweight U-Net with three resolution levels and two residual blocks per level, starting from 64 channels and doubling at each down-sample stage.
Generated Trajectories
$(a,c)=(1,1)$
$(a,c)=(25,5)$
$(a,c)=(100,10)$
$(a,c)=(225,15)$
$(a,c)=(900,30)$
Figure 11: Trajectories for $T=1$ at equidistant time points.
$(a,c)=(1,1)$
$(a,c)=(25,5)$
$(a,c)=(100,10)$
$(a,c)=(225,15)$
$(a,c)=(900,30)$
Figure 12: Trajectories for $T=5$ at equidistant time points.

### v2

@@ 5.2 Multi-Dimensional Kac Process
$\frac{C_{1}}{(4\pi\sigma^{2})^{\frac{d}{4}}}\left(\frac{1}{at^{1+\frac{d}{4}}}+\frac{1}{ca^{\frac{1}{2}}t^{\frac{1}{2}+\frac{d}{4}}}\right).$ (78)
@@
$\leq\Big{(}C_{1}\Big{(}\frac{1}{at}+\frac{1}{c(ta)^{\frac{1}{2}}}\Big{)}+\frac{1}{(4\pi\sigma^{2}t)^{\frac{1}{4}}}\Big{)}$ (81)
@@ 6 Velocity Field of the Multi-Dimensional Kac Flow
In this section, we want to develop a method how to retrieve the velocity field $v_{t}$ of the Kac flow in multiple dimensions. By conditioning on single points $x_{0}\in\mathbb{R}^{d}$, the multi-dimensional velocity field $v_{t}(x|x_{0})$ decomposes into univariate velocities $v_{t}^{i}(x^{i}|x_{0}^{i})$ which are analytically and numerically accessible. These conditional velocity fields can then be used to approximate the true velocity field $v_{t}$ within the framework of conditional flow matching. But note that conceptually, our velocities are conditioned to start at $t=0$ in the target measure – reversing the direction of usual flow matching.
@@
Furthermore, assume that the one-dimensional measure flows $\mu^{i}_{t}$ satisfy $\mu^{i}_{t}\in AC^{2}(I;\mathcal{P}_{2}(\mathbb{R}))$ and a continuity equation with some vector field $v^{i}_{t}$, i.e.
@@
By Proposition 10, $\mu_{t}(\cdot|x_{0})$ admits a velocity field $v_{t}(\cdot|x_{0})$ satisfying the continuity equation (2). By construction, the components $(X_{t}^{1},...,X_{t}^{d})$ of $\mathbf{X}_{t}$ are mutually independent. Hence, $\mu_{t}(x|x_{0})=\prod_{i=1}^{d}\mu_{t}^{i}(x^{i}|x_{0}^{i})$, where the one-dimensional probability flows $\mu_{t}^{i}(\cdot|x_{0}^{i})$ fulfill a continuity equation with 1D velocity fields $v_{t}^{i}(\cdot|x_{0}^{i})$. By Lemma 12, it holds the decomposition (86), and Theorem 4 immediately gives the formula (91). Finally, [19, Theorem 3.3], resp. [41, Theorem 6.3] provides (92). ∎
In stochastic analysis, the relation (92) has been known in connection with the mimicking of stochastic processes, see e.g. [3, 11]. Here, the mimicking process is given by the flow ODE (5).
Having identified the conditional vector field $v_{t}(\cdot|x_{0})$ for the conditional distribution $\mu_{t}(\cdot|x_{0})$ of the Kac process, we get the following result by employing [17], see also [41, Theorem 5.1, Remark 5.3], with slight adjustments to our setting.
@@
where $\mathcal{U}(0,T)\times_{t}u_{t}$ is understood in the sense of disintegration, see [41, Ch. 2.3], but here we have no direct access to $v$. However, following the same lines as in [17], we see that
@@
with a constant not depending on $\theta$, and the Conditional Flow Matching (CFM) loss is given by
@@
Here, we have access to both $v(t,x|x_{0})$ and $u_{t}(\cdot|x_{0})$ via (86) and (32), respectively, which we will exploit in the next section.
@@ 7 Numerical Experiments
In this section, we use the flow matching framework to train a neural network to approximate the velocity field from Corollary 14 by minimizing the CFM loss given in (95). To this end, we compute the conditional velocity $v_{t}(x|x_{0})$ via its analytic formula (91). In order to sample from $\mathbf{X}_{t}\sim u_{t}(\cdot|x_{0})$, where $x_{0}$ is fixed, we only need to sample i.i.d. components $X_{t}^{i}$ of the Kac process starting in 00. This can be done efficiently, since we have access to the analytic form (32) of the one-dimensional Kac distribution starting in 00 (use e.g. inverse transform sampling). Further, if we wish to sample fixed realizations of a Kac process for varying times $t$, we can efficiently simulate the random walk itself as explained in Appendix A.1.
During inference, we sample the latent variable from the Kac process $K_{T}$ starting in zero, at a final time $T$, and then integrate the flow ODE (5) backwards in time using an ODE solver with the learned vector field as the drift. Using the standard Kac flow introduces an approximation error, namely the distance between $X_{T}$ and $K_{T}$. This can be avoided by applying the mean reverting Kac flow (96) which we will introduce below.
7.1 Diffusion vs. Kac Process: A 2D Experiment with Dirac-like Modes
@@
We apply the conditional flow matching framework to both the Kac and diffusion based model. Here we use the Kac process (26) starting in the data with parameters $\frac{c^{2}}{a}=1$, and the standard diffusion with parameter $\sigma=1$ and the velocity field (11).
@@
7.2 Image Generation
In general, the ’standard’ Kac process $X_{t}=X_{0}+K_{t}$ and $K_{t}$ themselves are not equal at the final time $T$. It is common practice in generative modeling to augment the process so that the data no longer contributes to the final distribution (compare with variance-preserving diffusion processes, e.g. the Ornstein–Uhlenbeck process). Additionally, we introduce a time-scheduling function for the noising process. For simplicity, we fix $T=1$ in the rest of this section.
The Mean-Reverting Kac Process.
We define the mean-reverting Kac process by
$M_{t}\;\coloneqq\;f(t)\,X_{0}\;+\;K_{g(t)},\quad t\in[0,1],$ (96)
for smooth functions $f,g$ satisfying
$f(0)=1,\quad f(1)=0\quad\text{and}\quad g(0)=0,\quad g(1)=1.$
In particular, $M_{1}=K_{1}$. Differentiating (96) gives
$\dot{M}_{t}=\dot{f}(t)\,X_{0}\;+\;\dot{g}(t)\,\dot{K}_{g(t)},$
and hence the conditional velocity field is given by
$v_{M}(t,x\mid x_{0})$ $=\mathbb{E}\bigl{[}\dot{M}_{t}\mid M_{t}=x,\;X_{0}=x_{0}\bigr{]}$
$=\mathbb{E}\bigl{[}\dot{f}(t)\,x_{0}+\dot{g}(t)\,\dot{K}_{g(t)}\,\bigm{|}\,K_{g(t)}=x-f(t)x_{0},\;X_{0}=x_{0}\bigr{]}$
$=\dot{f}(t)\,x_{0}\;+\;\dot{g}(t)\,v_{K}\bigl{(}g(t),\,x-f(t)x_{0}\mid 0\bigr{)},$ (97)
where $v_{K}(t,x\mid 0)$ is the conditional Kac velocity field from (86). By the same arguments as in Corollary 14, one recovers the corresponding vector field for the mean-reverting Kac process.
Connection to Flow Matching.
Now let $R_{t}\coloneqq f(t)X_{0}+B_{g(t)}$ with a standard Brownian motion $B_{t}$, and choose
$f(t)=1-t,\qquad g(t)=t^{2}.$
Then it holds
[…]
Notice that (98) matches exactly the forward process used in recent flow matching works see [17, 41]. We experimented with various schedules and found that the above flow matching pair
$(f,g)=(1-t,\;t^{2})$
yielded good generation results. We compare our approach to those using the process $(R_{t})_{t}$, i.e. flow matching based on the independent coupling.
CIFAR-10.
We test our approach on generating CIFAR-10 images. We train the velocity field using conditional flow matching with the vector field (7.2) and compare a variety of choices for $(a,c)$. We apply a UNET architecture to parameterize the velocity field and use an ODE solver to simulate the corresponding flow ODE. We used the same setup employed in [39] to train our models. Our code is available online [^1] https://github.com/JChemseddine/telegraphers . For further details we refer to the implementation details in Appendix A.2.
In Table 1 we report the FID results for our methods as well as the baselines. Note that we implemented the baseline ourselves and show the resulting FID as "ours", and we show the reported result in [39] as "reported". In Figures 6, 7 we plot a selection of the generated images. Recall again that $a$ corresponds to the damping coefficient and $c$ to the propagation speed. We found that a combination of sufficiently high damping and lower speeds led to the best generation results. With this experiment we aim to demonstrate the scalability of our method, an extensive study of the effect of the parameters $(a,c)$ is not in the scope of the paper and left to future work.
Table 1: FID scores (100-step Euler sampling) for various $(a,c)$-Kac settings, plus the flow matching baseline.
Method FID Method FID
$a=900,\,c=10$ 18.08 $a=25,\,c=1$ 8.71
$a=900,\,c=20$ 43.39 $a=25,\,c=2$ 6.80
$a=900,\,c=30$ 67.97 $a=25,\,c=3$ 7.64
$a=100,\,c=1$ 12.12 $a=25,\,c=4$ 10.75
$a=100,\,c=3$ 10.78 $a=25,\,c=5$ 16.79
$a=100,\,c=5$ 17.31 FM (reported) 4.64
$a=100,\,c=7$ 26.97 FM (ours) 7.59
$a=100,\,c=10$ 41.34
$(a,c)=(25,2)$
$(a,c)=(25,3)$
FM
Figure 6: Generation results of the best performing mean-reverting Kac models and the baseline FM model.
$(a,c)=(100,1)$
$(a,c)=(100,3)$
$(a,c)=(900,10)$
Figure 7: Further generation results of the mean-reverting Kac models.
@@ Acknowledgement.
J. Chemseddine acknowledges funding of the DFG project STE 571/17-2 within the SPP 2298 "Theoretical Foundations of Deep Learning". R. Duong, P. Friz and G. Steidl acknowledge funding of the project EF-LI-Opt-2 within the DFG Excellence Cluster MATH+. The work of P. Friz is also supported by DFG CRC/TRR 388, project A02. We thank Gregor Kornhardt, Matthias Liero and Nicolaj Rux for fruitful discussions.
@@ A.2 Implementation Details
i.e. we draw $X_{0}$ directly from our GMM and add the exact Kac noise $K_{t}$, thereby avoiding any sampling approximation in this toy example. Analogously, in the diffusion case, we use a standard Brownian motion $B_{t}$ instead of $K_{t}$. All models were trained for $500{,}000$ iterations with a learning rate of $5\times 10^{-4}$ and a batch size of $256$. For validation we compute the negative log-likelihood (NLL) of $N$ model-generated samples under the true GMM density:
@@
CIFAR Experiment.
We trained all models for $400$k iterations, a batch size of $128$ and a learning rate of $2\times 10^{-4}$. We used the torchdiffeq [6] package to simulate the ODEs with an explicit Euler scheme. We use the time dependent U-Net architecture from [26]. We use the same network parameters as in [39], also employing an EMA with decay rate $0.9999$. Both training and evaluation was done using fixed random seeds. The FID is computed on $50$k samples using the torch fidelity package [29]. The code is available at https://github.com/JChemseddine/telegraphers.